# Testing quantum gravity effects with global lepton number symmetry and dark matter

### v1

@@ Abstract
We study the possibility of testing quantum gravity (QG) effects with a type-I seesaw origin of light neutrino mass embedded in a global $U(1)_{L}$ symmetric framework which also includes dark matter (DM). The spontaneous breaking of $U(1)_{L}$ symmetry generates the seesaw scale as well as DM mass dynamically. The (pseudo) Nambu-Goldstone boson, known as majoron, acquires non-zero mass due to soft symmetry breaking terms of quadratic type in the scalar potential, which eventually breaks $U(1)_{L}$ to its $Z_{2}$ subgroup. The spontaneous symmetry breaking, which effectively breaks $Z_{2}$, leads to the formation of domain walls (DW), posing a threat to successful cosmology, if allowed to dominate. As gravity does not respect any global symmetries, we consider higher dimensional operators suppressed by the scale of QG namely, $\Lambda_{\rm QG}$ which introduces the required bias leading to DW annihilation and emission of stochastic gravitational waves (GW) observable at near future experiments. The same operators also lead to decay of DM bringing interesting indirect detection aspects. While DM is produced non-thermally via scalar portal interactions, light majoron can give rise to additional $\Delta N_{\rm eff}$ within reach of future cosmic microwave background (CMB) experiments.
@@ I Introduction
In spite of having several motivations for BSM physics mentioned above, none of the experiments across different frontiers have observed any convincing signs of it so far. Canonical seesaw models consistent with neutrino mass and leptogenesis typically have a very high scale for new physics, away from direct experimental reach. On the other hand, WIMP DM has not shown up in direct search experiments either LUX-ZEPLIN:2022qhg . This might be indicative of the fact that DM is either feebly coupled or interact with the standard model (SM) particles via superheavy mediators. Such null results at DM search or other particle physics experiments searching for BSM physics have also motivated the search for complementary probes at other experiments. For example, some recent attempts have focused on probing high scale seesaw via stochastic gravitational wave (GW) observations Dror:2019syi ; Blasi:2020wpy ; Fornal:2020esl ; Samanta:2020cdk ; Barman:2022yos ; Huang:2022vkf ; Dasgupta:2022isg ; Okada:2018xdh ; Hasegawa:2019amx ; Borah:2022cdx ; Borah:2022vsu ; Barman:2023fad ; Borah:2023saq . The particle physics scenarios explored in these works considered the presence of additional symmetries whose spontaneous breaking not only lead to the dynamical generation of seesaw scale but also leads to the generation of stochastic GW from topological defects like cosmic strings Dror:2019syi ; Blasi:2020wpy ; Fornal:2020esl ; Samanta:2020cdk ; Borah:2022vsu , domain walls Barman:2022yos ; Barman:2023fad ; Saikawa:2017hiv ; Roshan:2024qnv ; Bhattacharya:2023kws or from bubbles generated at first order phase transition Huang:2022vkf ; Dasgupta:2022isg ; Okada:2018xdh ; Hasegawa:2019amx ; Borah:2022cdx ; Borah:2023saq . GW observations has also been used to probe even higher scales than seesaw, where quantum gravity (QG) effects can become important King:2023ayw ; King:2023ztb ; Borah:2022wdy . In these works, the QG effects on explicit global discrete symmetry breaking was considered in the context of domain wall formation from spontaneous discrete symmetry breaking and consequent GW emission. The authors of King:2023ayw ; King:2023ztb also considered the impact of such QG effects on decay of DM which is stabilised by discrete symmetry at renormalisable level.
@@
Motivated by these, we consider a global $U(1)$ symmetry related to the seesaw scale and DM stability and study QG effects leading to stochastic GW and DM decay. Connection to lepton sector and seesaw becomes natural if this Abelian global symmetry is identified as the global lepton number symmetry $U(1)_{L}$. In the simplest realisation of seesaw, the spontaneous breaking of $U(1)_{L}$ generates the right handed neutrino (RHN) masses or the seesaw scale dynamically. The same breaking also results in a massless Nambu-Goldstone (NG) boson, known as the majoron Chikashige:1980qk ; Chikashige:1980ui ; Gelmini:1980re ; Manna:2022gwn . Majoron can acquire a non-zero mass from explicit $U(1)_{L}$ breaking mass terms. When the singlet scalar responsible for breaking $U(1)_{L}$ spontaneously acquires a non-zero vacuum expectation value (VEV), it essentially breaks a $Z_{2}$ symmetry, remnant from $U(1)_{L}$ after explicit $U(1)_{L}$ breaking mass terms are included. Even without such explicit symmetry breaking terms, $U(1)_{L}$ is anomalous in the SM and broken to $Z_{3}$ by non-perturbative effects tHooft:1976rip ; Lazarides:2018aev  [^1] See Lazarides:2018aev ; Brune:2022vzd for non-minimal majoron models without domain walls.. Spontaneous breaking of discrete symmetries like $Z_{2}$ leads to the formation of domain walls (DW) Zeldovich:1974uw ; Kibble:1976sj ; Vilenkin:1981zs ; Saikawa:2017hiv ; Roshan:2024qnv , which can overclose the Universe, if allowed to dominate [^2] Spontaneous breaking of global $U(1)_{L}$ can also lead to formation of topological defects known as cosmic strings. However, GW emission remains suppressed compared to particle emission from global strings Baeza-Ballesteros:2023say .. We then consider higher dimensional operators suppressed by QG scale as bias leading to disappearance of DW thereby generating stochastic GW signal detectable at ongoing and future experiments. We consider a Dirac fermion having chiral transformation under $U(1)_{L}$ to be the DM which remains accidentally stable even after $U(1)_{L}$ or $Z_{2}$ breaking, unlike earlier works King:2023ayw ; King:2023ztb where a separate $Z_{2}$ symmetry was introduced for DM stability. The $U(1)_{L}$ global symmetry therefore, generates the scale of seesaw, DM mass, ensures the stability of DM and also leads to DW formation in the presence of $Z_{2}$-preserving explicit $U(1)_{L}$ breaking terms generating majoron mass. The higher dimensional operators suppressed by QG scale can however, lead to DM decay opening up some indirect detection prospects. We discuss the details of DM which interacts with the SM via singlet scalar portal, and the effects of QG scale suppressed operators on GW emission from DW and on DM phenomenology. The RHNs not only generate light neutrino masses via seesaw mechanism, but can also generate the baryon asymmetry of the Universe via leptogenesis Fukugita:1986hr . The emergence of $Z_{2}$ symmetry from a global $U(1)_{L}$, effects of QG on DW and DM decay, the presence of the pseudo NG boson or majoron bring complementary detection prospects at particle physics, gravitational waves as well as cosmology based experiments.
@@ II The Setup
The relevant field content and respective transformations under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{L}$ symmetry are shown in table 1 with the BSM fields being the right handed neutrino $N_{R}$, fermion DM $\chi$ and the singlet scalar $\Phi$ responsible for spontaneous symmetry breaking. The relevant part of the Yukawa Lagrangian can be written as
@@
which breaks $U(1)_{L}$ to a remnant $Z_{2}$ symmetry in the scalar potential.
@@
Finally, there exists no exact (continuous or discrete) global symmetry in the theory of QG Kallosh:1995hi . In other words, any global symmetry of a given effective field theory (EFT) is at best an approximate symmetry emergent in the IR Witten:2017hdv and should be broken by a higher-dimensional operator. As a result, the remnant $Z_{2}$ present in this setup will also be explicitly broken by the higher dimensional operators Rai:1992xw ,
@@ III Cosmology of majoron and dark matter
Here $\rho_{\eta}$ and $\rho_{\nu,1}$ denotes the energy density of majoron and a single species of active neutrino respectively and $g_{*}$ denotes the effective degrees of freedom of SM particles. For all the blue shaded region shown in the right panel plot of Fig. 1, majoron gets decoupled at a temperature larger than $100$ GeV. This gives a constant contribution of thermal majoron $\Delta N_{\rm eff}=0.027$ for all $v_{\Phi}$. Future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ upto $0.014$ at $1\sigma$, keeping thermalised light majoron parameter space verifiable.
@@ III.3 Leptogenesis
Here $(\Delta m_{\rm atm})^{2}\simeq 2.4\times 10^{-3}\rm eV^{2}$ denotes the active neutrino atmospheric mass-squared difference. We consider efficiency factor to be $0.01$ that corresponds to strong wash-out regime. Apart from usual leptogenesis in type-I seesaw, in this setup we have extra annihilation channel, $N\bar{N}\to\eta\eta$ that can keep RHNs in equilibrium for longer and hence dilute the final baryon asymmetry. However for strong washout regime and for equilibrium initial abundance of $N$, the effect from $N\bar{N}\to\eta\eta$ can be neglected. With this, using Eq. (12), we get that the mass of RHN should be larger than $m_{N_{1}}\gtrsim 6\times 10^{10}$ GeV. As the DM analysis is valid for $m_{\phi}>2m_{N}$ as discussed in the previous section, this requires $v_{\Phi}$ to be larger than $\sim 3\times 10^{12}$ GeV for $\lambda_{\Phi}=10^{-3}$. Such large values of $v_{\Phi}$ are disfavoured from DW overclosure and DM properties, as we discuss below. For a $v_{\Phi}$ smaller than $3\times 10^{12}$ GeV, required amount of lepton asymmetry can be created via resonant leptogenesis Pilaftsis:2003gt . Resonant leptogenesis can enhance CP asymmetry parameter upto $\mathcal{O}(0.1)$ where two RHNs have nearly the same mass, with the mass splitting being nearly equal to the its decay width. Therefore, correct baryon asymmetry is possible via resonant scenario where $v_{\Phi}$ is as small as $10^{6}$ GeV.
@@ V Domain walls and Gravitational Waves
Domain walls, a cosmological catastrophe, are 2-dimensional topological defects resulting from a spontaneous symmetry breaking (SSB) of a discrete symmetry like $Z_{2}$, as also considered in the present setup. Once formed, their energy density varies inversely with the scale factor ($a^{-1}$) which is much slower than that of matter ($a^{-3}$) or radiation ($a^{-4}$). As a result, they can dominate the energy budget of the Universe at some later stage and can alter the fate of CMB observations or successful light nuclei synthesis during BBN. These cosmological catastrophes can be avoided by introducing an energy bias in the scalar potential that makes the vacuum unstable. In the present setup this can be achieved with the help of QG effect as it is expected to break any global symmetry at high scale. Following this, one can write the energy bias after replacing fields with their VEVs, this leads to a bias contribution to the effective potential of the form
@@
Due to this uplifting of degeneracy, the DW start to collapse and annihilate, resulting in the production of SGWB whose signals can be tested in the present and future GW detectors. Under the assumption that $v_{\Phi}\gg v$, a sufficiently large $V_{\rm bias}$ can be achieved rendering the production of observable GW signals. Note that to have a DW, a very large $V_{\rm bias}$ must be avoided according to the prediction of percolation theory Saikawa:2017hiv . Under such a hierarchy, the bias potential is dominated by,
@@
where $C_{\rm ann}\simeq 2$ is a dimensionless constant, $\mathcal{A}\simeq 0.8$ denotes the area parameter and $\sigma=\sqrt{\frac{8\lambda_{\Phi}}{9}}v_{\Phi}^{3}$ is the surface tension of the DW. Even if Eq. (21) is satisfied and the DW are annihilated away before their dominance, their decay products (assuming they are the SM particles) can still destroy the light elements created at BBN epoch. This demands $t_{\rm ann}\leq 0.01$ s and hence an additional constraint on the magnitude of the energy bias can be obtained as,

### v2

# Observable gravitational waves and $\Delta N_{\rm eff}$ with global lepton number symmetry and dark matter

@@ Abstract
We study the possibility of testing a dark matter (DM) scenario embedded in a global lepton number symmetry $U(1)_{L}$ via gravitational waves (GW) and cosmic microwave background (CMB) observations. The spontaneous breaking of $U(1)_{L}$ symmetry generates the seesaw scale as well as DM mass dynamically. The (pseudo) Nambu-Goldstone boson, known as majoron, acquires non-zero mass due to soft symmetry breaking terms of quadratic type in the scalar potential, which eventually breaks $U(1)_{L}$ to its $Z_{2}$ subgroup. The spontaneous symmetry breaking, which effectively breaks $Z_{2}$, leads to the formation of domain walls (DW), posing a threat to successful cosmology, if allowed to dominate. As gravity does not respect any global symmetries, we consider higher dimensional operators suppressed by the scale of quantum gravity (QG) namely, $\Lambda_{\rm QG}$ which introduces the required bias leading to DW annihilation and emission of stochastic gravitational waves (GW) observable at near future experiments. The same operators also lead to decay of DM bringing interesting indirect detection aspects. While DM is produced non-thermally via scalar portal interactions, light majoron can give rise to additional $\Delta N_{\rm eff}$ within reach of future CMB experiments.
@@ I Introduction
In spite of having several motivations for BSM physics mentioned above, none of the experiments across different frontiers have observed any convincing signs of it so far. Canonical seesaw models consistent with neutrino mass and leptogenesis typically have a very high scale for new physics, away from direct experimental reach. On the other hand, WIMP DM has not shown up in direct search experiments either LUX-ZEPLIN:2022qhg . This might be indicative of the fact that DM is either feebly coupled or interact with the standard model (SM) particles via superheavy mediators. Such null results at DM search or other particle physics experiments searching for BSM physics have also motivated the search for complementary probes at other experiments. For example, some recent attempts have focused on probing high scale seesaw via stochastic gravitational wave (GW) observations Dror:2019syi ; Blasi:2020wpy ; Fornal:2020esl ; Samanta:2020cdk ; Barman:2022yos ; Huang:2022vkf ; Dasgupta:2022isg ; Okada:2018xdh ; Hasegawa:2019amx ; Borah:2022cdx ; Borah:2022vsu ; Barman:2023fad ; Borah:2023saq . The particle physics scenarios explored in these works considered the presence of additional symmetries whose spontaneous breaking not only lead to the dynamical generation of seesaw scale but also leads to the generation of stochastic GW from topological defects like cosmic strings Dror:2019syi ; Blasi:2020wpy ; Fornal:2020esl ; Samanta:2020cdk ; Borah:2022vsu , domain walls Barman:2022yos ; Barman:2023fad ; Saikawa:2017hiv ; Roshan:2024qnv ; Bhattacharya:2023kws ; Blasi:2022ayo ; Blasi:2023sej or from bubbles generated at first order phase transition Huang:2022vkf ; Dasgupta:2022isg ; Okada:2018xdh ; Hasegawa:2019amx ; Borah:2022cdx ; Borah:2023saq . GW observations has also been used to probe even higher scales than seesaw, where quantum gravity (QG) effects can become important King:2023ayw ; King:2023ztb ; Borah:2022wdy . In these works, the QG effects on explicit global discrete symmetry breaking was considered in the context of domain wall formation from spontaneous discrete symmetry breaking and consequent GW emission. The authors of King:2023ayw ; King:2023ztb also considered the impact of such QG effects on decay of DM which is stabilised by discrete symmetry at renormalisable level.
@@
Motivated by these, we consider a global $U(1)$ symmetry related to the seesaw scale and DM stability and study QG effects leading to stochastic GW and DM decay. Connection to lepton sector and seesaw becomes natural if this Abelian global symmetry is identified as the global lepton number symmetry $U(1)_{L}$. In the simplest realisation of seesaw, the spontaneous breaking of $U(1)_{L}$ generates the right handed neutrino (RHN) masses or the seesaw scale dynamically. The same breaking also results in a massless Nambu-Goldstone (NG) boson, known as the majoron Chikashige:1980qk ; Chikashige:1980ui ; Gelmini:1980re ; Manna:2022gwn . Majoron can acquire a non-zero mass from explicit $U(1)_{L}$ breaking mass terms. When the singlet scalar responsible for breaking $U(1)_{L}$ spontaneously acquires a non-zero vacuum expectation value (VEV), it essentially breaks a $Z_{2}$ symmetry, remnant from $U(1)_{L}$ after explicit $U(1)_{L}$ breaking mass terms are included. Even without such explicit symmetry breaking terms, $U(1)_{L}$ is anomalous in the SM and broken to $Z_{3}$ by non-perturbative effects tHooft:1976rip ; Lazarides:2018aev  [^1] See Lazarides:2018aev ; Brune:2022vzd for non-minimal majoron models without domain walls.. Spontaneous breaking of discrete symmetries like $Z_{2}$ leads to the formation of domain walls (DW) Zeldovich:1974uw ; Kibble:1976sj ; Vilenkin:1981zs ; Saikawa:2017hiv ; Roshan:2024qnv , which can overclose the Universe, if allowed to dominate [^2] Spontaneous breaking of global $U(1)_{L}$ can also lead to formation of topological defects known as cosmic strings Fu:2023nrn . However, GW emission remains suppressed compared to particle emission from global strings Baeza-Ballesteros:2023say .. We then consider higher dimensional operators suppressed by QG scale as bias leading to disappearance of DW thereby generating stochastic GW signal detectable at ongoing and future experiments. We consider a Dirac fermion having chiral transformation under $U(1)_{L}$ to be the DM which remains accidentally stable even after $U(1)_{L}$ or $Z_{2}$ breaking, unlike earlier works King:2023ayw ; King:2023ztb where a separate $Z_{2}$ symmetry was introduced for DM stability. The $U(1)_{L}$ global symmetry therefore, generates the scale of seesaw, DM mass, ensures the stability of DM and also leads to DW formation in the presence of $Z_{2}$-preserving explicit $U(1)_{L}$ breaking terms generating majoron mass. The higher dimensional operators suppressed by QG scale can however, lead to DM decay opening up some indirect detection prospects. We discuss the details of DM which interacts with the SM via singlet scalar portal, and the effects of QG scale suppressed operators on GW emission from DW and on DM phenomenology. The RHNs not only generate light neutrino masses via seesaw mechanism, but can also generate the baryon asymmetry of the Universe via leptogenesis Fukugita:1986hr . The emergence of $Z_{2}$ symmetry from a global $U(1)_{L}$, effects of QG on DW and DM decay, the presence of the pseudo NG boson or majoron bring complementary detection prospects at particle physics, gravitational waves as well as cosmology based experiments.
@@ II The Setup
The relevant field content and respective transformations under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{L}$ symmetry are shown in table 1 with the BSM fields being the right handed neutrino $N_{R}$, fermion DM $\chi$ and the singlet scalar $\Phi$ responsible for spontaneous symmetry breaking. While the specific choice of $Q_{L}$, the lepton number charge of $\chi_{L}$ does not affect our conclusion, we consider $Q_{L}\neq-1$, such that DM remains stable at renormalisable level. The relevant part of the Yukawa Lagrangian can be written as
@@
which breaks $U(1)_{L}$ to a remnant $Z_{2}$ symmetry in the scalar potential. It should be noted that we have considered the simplest form of explicit $U(1)_{L}$ breaking to generate the mass of majoron. Other forms of explicit $U(1)_{L}$ breaking will lead to different remnant symmetry. For example, explicit $U(1)_{L}$ breaking terms like $\Phi^{3},\Phi^{4}$ will correspond to remnant symmetry $Z_{3},Z_{4}$ respectively. Allowing, the explicit breaking to occur via non-renormalisable operators $\Phi^{n}/\Lambda^{n-4}$ will lead to bigger remnant symmetry groups like $Z_{n}$ where $n>4$. The details of the domain wall formation, evolution and emission of GW differs from the simplest $Z_{2}$ domain walls discussed in this work. Phenomenology of such $Z_{n}\,(n>2)$ domain walls can be found in Wu:2022stu ; Wu:2022tpe ; Bai:2023cqj and references therein.
@@
Finally, there exists no exact (continuous or discrete) global symmetry in the theory of QG Kallosh:1995hi . In other words, any global symmetry of a given effective field theory (EFT) is at best an approximate symmetry emergent in the IR Witten:2017hdv and should be broken by a higher-dimensional operator. As a result, the remnant $Z_{2}$ present in this setup will also be explicitly broken by the higher dimensional operators Rai:1992xw  [^3] See Beyer:2022ywc and references therein for similar discussions in QCD axion models with Peccei-Quinn $U(1)$ global symmetry.,
@@ III Cosmology of majoron and dark matter
Here $\rho_{\eta}$ and $\rho_{\nu,1}$ denotes the energy density of majoron and a single species of active neutrino respectively and $g_{*}$ denotes the effective degrees of freedom of SM particles. For all the blue shaded region shown in the right panel plot of Fig. 1, majoron gets decoupled at a temperature larger than $100$ GeV. This gives a constant contribution of thermal majoron $\Delta N_{\rm eff}=0.027$ for all $v_{\Phi}$. Future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ upto $0.014$ at $1\sigma$, keeping thermalised light majoron parameter space verifiable. We also check the non-thermal production of majorons from topological defects like cosmic strings and domain walls in our model Reig:2019sok and find the corresponding abundance to be sub-dominant compared to DM, for the region of parameter space considered in our work.
@@ III.3 Leptogenesis
Here $(\Delta m_{\rm atm})^{2}\simeq 2.4\times 10^{-3}\rm eV^{2}$ denotes the active neutrino atmospheric mass-squared difference. We consider efficiency factor to be $0.01$ that corresponds to strong wash-out regime. Apart from usual leptogenesis in type-I seesaw, in this setup we have extra annihilation channel, $N\bar{N}\to\eta\eta$ that can keep RHNs in equilibrium for longer and hence dilute the final baryon asymmetry. However for strong washout regime and for equilibrium initial abundance of $N$, the effect from $N\bar{N}\to\eta\eta$ can be neglected. With this, using Eq. (12), we get that the mass of RHN should be larger than $m_{N_{1}}\gtrsim 6\times 10^{10}$ GeV. Imposing the condition of correct DM abundance along with correct baryon asymmetry, one needs $m_{\phi}>2m_{N}$ $\implies v_{\Phi}>\frac{2m_{N}}{\sqrt{2\lambda_{\Phi}}}$. To make the calculations consistent with the DM and GW analysis, we consider $\lambda_{\Phi}=10^{-3}$ which gives $v_{\Phi}$ larger than $\sim 3\times 10^{12}$ GeV. Such large values of $v_{\Phi}$ are disfavoured from DW overclosure and DM properties, as we discuss below. With the help of flavor leptogenesis, the minimum $m_{N_{1}}$ can be relaxed upto one order of magnitude Blanchet:2006be ; Dev:2017trv than the value obtained from Eq. (12). Hence the corresponding value of $v_{\Phi}$ can also be reduced upto one order magnitude. For a even smaller value of $v_{\Phi}$, required amount of lepton asymmetry can be created via resonant leptogenesis Pilaftsis:2003gt . Resonant leptogenesis can enhance CP asymmetry parameter upto $\mathcal{O}(0.1)$ where two RHNs have nearly the same mass, with the mass splitting being nearly equal to the its decay width. Therefore, correct baryon asymmetry is possible via resonant scenario where $v_{\Phi}$ is as small as $10^{6}$ GeV.
@@ V Domain walls and Gravitational Waves
Domain walls, a cosmological catastrophe, are 2-dimensional topological defects resulting from a spontaneous symmetry breaking (SSB) of a discrete symmetry like $Z_{2}$, as also considered in the present setup. Once formed, their energy density varies inversely with the scale factor ($a^{-1}$) which is much slower than that of matter ($a^{-3}$) or radiation ($a^{-4}$). As a result, they can dominate the energy budget of the Universe at some later stage and can alter the fate of CMB observations or successful light nuclei synthesis during BBN. These cosmological catastrophes can be avoided by introducing an energy bias in the scalar potential that makes the vacuum unstable. The above scaling of the DW is true assuming the radiation-dominated era when the DWs were relativistic. At a later stage, the DW dynamic is dominated by its tension force that stretches the DWs up to the horizon size. Numerical studies Press:1989yh ; Garagounis:2002kt ; Oliveira:2004he ; Leite:2011sc ; Leite:2012vn have shown that the evolution of DW in this regime can be described by the scaling solution. Here, their energy density evolves according to the simple scaling law $\rho_{\rm DW}\propto t^{-1}$ where $\rho_{\rm DW}$ denotes the energy density of the DW and their typical size is given by Hubble radius $\sim t^{2}$. In the scaling regime, their energy density can be expressed as,
$\rho_{\rm DW}=\sigma\frac{\mathcal{A}}{t},$ (19)
with $\mathcal{A}\simeq 0.8$ being the area parameter Hiramatsu:2013qaa and $\sigma$ being the surface tension of the DW. In the present setup this can be achieved with the help of QG effect as it is expected to break any global symmetry at high scale. Following this, one can write the energy bias after replacing fields with their VEVs, this leads to a bias contribution to the effective potential of the form
@@
Due to this uplifting of degeneracy, the DW starts to collapse and annihilate, resulting in the production of SGWB whose signals can be tested in the present and future GW detectors. Under the assumption that $v_{\Phi}\gg v$, a sufficiently large $V_{\rm bias}$ can be achieved rendering the production of observable GW signals. Note that to have a DW, a very large $V_{\rm bias}$ must be avoided according to the prediction of percolation theory Saikawa:2017hiv . Under such a hierarchy, the bias potential is dominated by,
@@
where $C_{\rm ann}\simeq 2$ is a dimensionless constant, and $\sigma=\sqrt{\frac{8\lambda_{\Phi}}{9}}v_{\Phi}^{3}$ is the surface tension of the DW. Even if Eq. (22) is satisfied and the DW are annihilated away before their dominance, their decay products (assuming they are the SM particles) can still destroy the light elements created at BBN epoch. This demands $t_{\rm ann}\leq 0.01$ s and hence an additional constraint on the magnitude of the energy bias can be obtained as,